2. A manufacturer of yoga pants sells them for $28 each. They hired some consultants who determined that the cost of manufacturing x pants was
C( x ) = x^2 − 2x − 9.
a)Write a function for the revenue (the amount of money the company brings in).
b)Write a function for the profit (the revenue - cost).
c)Find the number of t-shirts they should make to maximize the profit function. Round your answer to the nearest whole number.

Answer by ikleyn(52781)

.
2. A manufacturer of yoga pants sells them for $28 each. They hired some consultants who determined
that the cost of manufacturing x pants was C( x ) = x^2 − 2x − 9.
a)Write a function for the revenue (the amount of money the company brings in).
b)Write a function for the profit (the revenue - cost).
c)Find the number of t-shirts they should make to maximize the profit function. Round your answer to the nearest whole number.
~~~~~~~~~~~~~~~~~~


(a)  The revenue = 28*x dollars.



(b)  Profit = Revenue - Cost = 28x - (x^2 - 2x -9) = -x^2 + 30x + 9.



(c)  This question, as it is formulated and presented, is not relevant to the given part.


-----------

Do you really control on what you are writing ?